Flashlight batteries are sold at Batteries Are us in two-packs and four-packs. A Batteries Are Us clerk hangs six packages of them on each of twelve hooks. He hung 200 batteries in all. How many two-packs are in stock?
solve a system of equations
i'm confused with the 2-pack vs 4 pack scenario.
i need to show all work
please help!

Answer by stanbon(75887)

Flashlight batteries are sold at Batteries Are us in two-packs and four-packs. A Batteries Are Us clerk hangs six packages of them on each of twelve hooks. He hung 200 batteries in all. How many two-packs are in stock?
solve a system of equations
-------
Packs Equation:: t + f = 12*6 = 72 packs
Barreries Eq::: 2t + 4f = 200 batteries
----------------------------------------
Modify for elimination:
2t + 2f = 144
2t + 4f = 200
-----
Subtract and solve for "f":
2f = 56
f = 28 (# of 4-battery packs)
---
Solve for "t":
t + f = 72
28 + f = 72
f = 44 (# of 2-battery packs)
---------------------------------
Cheers,
Stan H.
=====================